As a logging truck rounds a bend in the road, some logs come loose and begin to roll without slipping down the mountainside. The mountain slopes upward at 48.0 degrees above the horizontal, and we can model the logs as solid uniform cylinders of mass 575 kg and diameter 1.20 m. Find the acceleration of the logs as they roll down the mountain.
Answer #1

acceleration of a cylinder rolling down = - g sinθ / ( 1 +Icom/MR2 )
for uniform cylinder Icom = 0.5MR2
∴ acceleration = - 9.8*sin48 / ( 1 + 0.5)
= - 4.855 m/s2 ( in the negativedirection of the x - axis )
Answer #2

Given that,
Angle between the horizontal and road θ=

Mass of the cylinder M=575 kg

Diameter of the cylinder d =1.20 m

then radius of cylinder 

a) Using parallel axis theorem,

the moment of inertia about the point of contact of the log with slope is 

the torque about the point of contact is τ=MgR sinθ

But we know that torque is τ=Iα

∴Iα=MgR sinθ

Angular acceleration α=

therefore acceleration a=αR

                                   a= (2gsin θ)/3

                                      =(2*9.8*sin 48)/3

                                      =4.8552

b) Since friction force provides the torque  to produce the angular acceleration about the center of the  mass

here f= frictional force

Moment of inertia about the center of the mass 

f=[( 575)*(9.8)*sin 48]/3=1395.8 N

c) it is a static friction
Answer #3

a) acceleration = - g sinθ / ( 1 +Icom/MR2 )
for uniform cylinder Icom =0.5MR2
∴ acceleration = - 9.8*sin48 / ( 1 + 0.5) = - 4.855 m/s2 ( in the negativedirection of the x - axis )


b) frictional froce is static.
∴ fs = - Icom a/ R2


=- 0.5Ma


= 1395N